OpenStudy (anonymous):

Point J is is located at (4,2) and point K is (12,-2) find the midpoint of segment JK.

OpenStudy (anonymous):

Alright, the midpoint can be found by adding the x coordinates together and dividing by two and the the same can be done to the y coordinates. For instance, (4+12)/2 = 8=x (2+-2)/2 = 0/2 = 0 =y Therefore (x,y) is the midpoint, which is (8,0). Does that make sense?

OpenStudy (anonymous):

yes! thank you !